OpenStudy (anonymous):

What is 3 1/2 divided by 2? Please show steps.

OpenStudy (anonymous):

Type in your calculator \[3.5\div2\] and you will get 1.75. Does this help??

OpenStudy (anonymous):

3 1/2 divided by 2= Look at 3 1/2 3 + 1/2 divided by 2 So 3 x 2 +1 divided by 2 = 6+1 divided by 2 = 7/2 Using this formula. A/B divided by c= A/ BxC 7/2x2= 7/4 7 divided by 4= 1.75